OpenStudy (anonymous):

Explain how to solve the equation 3t + 11 = –10. Show that your solution works.

OpenStudy (nenadmatematika):

3t=-10-11 3t=-21 t=-7 if you plug back into the equation this value for t you will get -10=-10 which means that you're right :D

OpenStudy (anonymous):

thanks

OpenStudy (nenadmatematika):

you're welcome :D

OpenStudy (anonymous):

3t=-10-11 (by subtracting both side by -11) 3t=-21 (divide by 3 both sides) t=-7

OpenStudy (anonymous):

thanks:)